HUBBLE SPACE TELESCOPE DAILY REPORT #2827 PERIOD COVERED: 0000Z (UTC) 03/12/01 - 0000Z (UTC) 03/13/01 Daily Status Report as of 072/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8581 (A Search For Low-Mass Companions To Ultracool Dwarfs) The WF/PC-2 was used to search for very low-mass {VLM} companions to a complete sample of 120 late-M and L dwarfs, drawn mainly from the 2MASS and SDSS surveys. The primary goal is to determine the multiplicity of M < 0.1 M_odot dwarfs. In particular, we aim to identify binary systems suitable for long-term astrometric monitoring and mass measurement, and systems with cool, sub-1000K companions. The proposal completed with no reported problems. 1.2 Completed Four Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.3 Completed WF/PC-2 8816 (Cycle 9 UV Earthflats) The WF/PC-2 was used to obtain sequences of Earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set and in order to monitor flat field stability. There were no reported problems. 1.4 Completed Two Sets of STIS/CCD 8865 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.5 Completed Two Sets of WF/PC-2 8583 (Imaging Snapshots of Asteroids) The WF/PC-2 was used to obtain images of the fifty largest main belt asteroids that have favorable apparitions during cycle 9. The images will be searched for companion bodies, as well as mineralogical variegation on the resolved main bodies. There were no reported problems 1.6 Completed Six Sets of WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed nominally. 1.7 Completed Two Sets of STIS/CCD 8591 (The Smallest Nuclear Black Holes) The Space Telescope Imaging Spectrograph (CCD) was used to observe small nuclear black holes which are the last major unexplored part of BH parameter space, searching for the smallest BHs that HST can possibly find. The proposal completed with no reported anomalies. 1.8 Completed Three Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The proposal completed with no reported problems. 1.9 Completed Two Sets of WF/PC-2 8817 (Cycle 9 INTFLAT Sweeps and Linearity Test) The WF/PC-2 was used to obtain INTFLAT observations in order to monitor the pixel to pixel flatfield response and provide a linearity check. The INTFLAT sequences, to be done once during the year, are similar to those from the Cycle 8 program 8448. The images will provide a backup database in the event of complete failure of the VISFLAT lamp as well as allow monitoring of the gain ratios. The sweep is a complete set of internal flats, cycling through both shutter blades and both gains. The linearity test consists of a series of INTFLATs in F555W, in each gain and each shutter. There was no anomalous activity. 1.10 Completed Six Sets of WF/PC-2 8811 (Cycle 9 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.11 Completed WF/PC-2 8812 (Cycle 9 Internal Monitor) The WF/PC-2 was used to obtain routine internal monitors for WFPC2 in order to monitor the health of the cameras. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. There were no reported problems. 1.12 Completed Two Sets of STIS/MA1/MA2 8843 (Cycle 9 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform the routine monitoring of the MAMA detector dark noise. The proposal completed nominally. 1.13 Completed STIS/CCD 8623 (HST, Chandra, and FUSE Studies of Interstellar Material toward HD 24534 {aka X Persei}) The Space Telescope Imaging Spectrograph (CCD) was used to make high-resolution UV spectroscopic observations of the reddened star HD 24534 {aka X Persei}, in order to: {1} establish accurate gas-phase column densities for elements other than carbon and oxygen {for which high- quality data already exist}, for the purpose of deriving {in conjunction with FUSE and Chandra observations} absolute depletions of several key elements; and {2} further pursue molecular abundances toward this star, to supplement HST/GHRS data obtained earlier which showed that this line of sight supports a rich chemistry. The observations completed nominally. 1.14 Completed WF/PC-2 8894 (Accretion Disks in the Orion Nebula) The WF/PC-2 was used to perform UB-band imaging of the Orion Nebula Cluster in order to derive, for the first time, the disk accretion rates of the pre-main-sequence objects, and to study with, unprecedented resolution and sensitivity the dark silhouette disks. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 3 Successful: 3 Scheduled Re-acquisitions: 11 Successful: 11 2.2 FHST Updates: Scheduled: 11 Successful: 11 2.3 Operations Notes: The SSR EDAC error counter was cleared four times per ROP SR-1A. An operations request was utilized to select the hardware safemode recorder at 071/2354Z. The ephemeris tables were uplinked at 072/0041Z as directed by ROP DF-07A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW